# Novel Dynamical Systems with Finite-Time and Fixed-Time Stability for Generalized Inverse Mixed Variational Inequality Problems

### v1

@@ Abstract
Two novel continuous-time dynamical systems are proposed to analyze the finite-time and fixed-time stability of solutions to GIMVIPs in finite-dimensional Hilbert spaces. Under suitable assumptions on the operators and model parameters, Lyapunov-based techniques are employed to establish finite-time and fixed-time convergence of the generated trajectories.
While both systems exhibit accelerated convergence, the settling time of the finite-time stable system depends on the initial condition, whereas the fixed-time stable system admits a uniform upper bound on the convergence time that is independent of the initial state. Moreover, an explicit forward Euler discretization of the continuous-time dynamics leads to a proximal point-type algorithm that preserves the fixed-time convergence property. Rigorous convergence analysis of the resulting iterative scheme is provided. A numerical experiment is presented to demonstrate the effectiveness of the proposed methods.
@@
Finite-time stability, Fixed-time Stability, proximal point Method, Dynamical System, generalized Inverse mixed Variational Inequality Problem.
@@ 1 Introduction
Let $\langle\cdot,\cdot\rangle$ and $\|\cdot\|$ denote the inner product and the associated norm in $\mathbb{R}^{d}$, respectively. This paper is concerned with a class of inverse equilibrium problems known as the generalized inverse mixed variational inequality problem (GIMVIP). Given a nonempty closed and convex set $\Omega\subset\mathbb{R}^{d}$, the GIMVIP consists in finding a vector $\overline{w}\in\mathbb{R}^{d}$ such that
$F(\bar{w})\in\Omega\quad\text{and}\quad\langle h(\bar{w}),v-F(\bar{w})\rangle+g(v)-g(F(\bar{w}))\geq 0,\quad\forall v\in\Omega,$ (1)
@@
The formulation (1) unifies and generalizes several well-known models in variational analysis. If $h$ is the identity mapping, then (1) reduces to the inverse mixed variational inequality problem introduced in [17]. When $g\equiv 0$, the GIMVIP collapses to the inverse variational inequality problem (IVIP), originally proposed by He and Liu [12]. On the other hand, if $F$ is the identity operator, then (1) becomes a mixed variational inequality problem and further reduces to the classical variational inequality problem (VIP) when $g\equiv 0$.
@@
Variational inequality problems constitute a fundamental framework for modeling equilibrium phenomena arising in optimization, saddle-point problems, Nash equilibria in noncooperative games, and fixed-point theory; see, for example, [3, 4]. Beyond direct formulations, inverse variational inequality problems have attracted increasing attention due to their capability to infer unknown system parameters from observed equilibrium states. Such inverse models naturally arise in applications where equilibrium behavior is observable, whereas the underlying cost structures, control parameters, or governing operators are not directly accessible.
@@
In recent years, inverse variational inequality models have been successfully applied to traffic network analysis, economic equilibrium, and telecommunication systems. To enhance modeling flexibility, inverse mixed variational inequality problems were proposed by incorporating nonsmooth and composite terms, thereby allowing for richer representations of practical constraints. Further developments have led to generalized inverse mixed variational inequality frameworks, which subsume many existing inverse models and provide a unifying setting for inverse equilibrium analysis. Motivated by these advances, the present work focuses on the GIMVIP as a versatile and comprehensive inverse equilibrium model.
@@
Alongside model generalization, the design of efficient solution methods remains a central challenge. Continuous-time dynamical systems have emerged as a powerful tool for solving variational inequalities and related equilibrium problems. By reformulating equilibrium conditions as dynamical processes, one can exploit Lyapunov stability theory to analyze convergence behavior. Compared with purely discrete-time schemes, continuous-time approaches offer conceptual simplicity, real-time implementation potential, and a natural feedback control interpretation.
@@
Most existing dynamical approaches for variational and inverse variational inequality problems guarantee asymptotic or exponential stability. More specifically, dynamical system techniques have been applied to optimization problems [1, 5, 10, 11, 18, 21], sparse optimization [9, 11, 26], control problems including multi-agent control systems [27], mixed variational inequality problems (MVIPs) [8, 13], variational inequality problems (VIPs), and equilibrium problems (EPs) [14, 15]. More recently, Nam et al. [24] investigated finite-time stability for generalized monotone inclusion problems, while fixed-time stable dynamical systems for such inclusion models were developed in [25]. In addition, Ju et al. [16] employed dynamical systems to study fixed-time stability for pseudo-monotone mixed variational inequality problems.
@@
As for inverse quasi-variational inequality problems, several works have employed dynamical systems to analyze their solution behavior; see, for instance, [28]. In 2023, Dey and Reich [7] proposed a first-order continuous-time dynamical system for inverse quasi-variational inequality problems and established global exponential stability. Subsequently, Vuong and Thanh [23] analyzed a time-discretized version of a similar dynamical system and obtained linear convergence results under a more general moving set condition $\Phi(w)$. More recently, Tangkhawiwetkul [22] employed dynamical system techniques to study generalized inverse mixed variational inequality problems and derived asymptotic as well as exponential stability results.
@@
However, in many applications requiring rapid and predictable convergence, asymptotic or exponential stability may be insufficient. The notion of finite-time stability, introduced in [2], ensures convergence within a finite settling time, although this time generally depends on the initial state. To overcome this drawback, Polyakov [20] introduced the concept of fixed-time stability, for which the convergence time admits a uniform upper bound independent of the initial condition. Despite their importance, finite-time and fixed-time convergence properties have not yet been systematically investigated for generalized inverse mixed variational inequality problems.
@@
Motivated by this gap, in this paper, we develop novel first-order continuous-time dynamical systems for solving GIMVIPs and establish their finite-time and fixed-time stability properties. By applying a forward Euler discretization, we further derive a proximal point-type iterative algorithm and show that the generated sequence converges to the unique solution of the GIMVIP within a fixed time. The proposed framework not only extends existing dynamical approaches for inverse variational inequality problems but also provides an efficient and robust method for achieving fast convergence in generalized inverse equilibrium models.
The paper is structured as follows. Section 2 introduces the necessary background material, including the concepts of finite-time and fixed-time stability. A proximal dynamical system related to GIMVIPs is then formulated, and the correspondence between its equilibrium points and the solutions of GIMVIPs is established. Several auxiliary technical results are also provided. In Section 3, a new dynamical system is developed, and its finite-time convergence properties are rigorously analyzed. Section 4 addresses the fixed-time stability of an alternative dynamical system tailored to GIMVIPs. Section 5 examines a sufficient condition for consistent discretization, ensuring fixed-time convergence of the discrete-time scheme within a bounded number of iterations. Section 6 presents a numerical simulation to validate the performance and advantages of the proposed discretization-based algorithm. Finally, Section 7 summarizes the main findings and outlines concluding remarks.
@@ 2.3 A nominal dynamical system
    $\langle w-\overline{w},\Xi(w)\rangle\geq(\rho-\Lambda)\|w-\overline{w}\|^{2}>0\ \forall w\in\mathbb{R}^{d},\ w\neq\overline{w}$ for all $w\in R^{d}$.
@@ 3 Finite-time stability analysis
4 Fixed-time stability analysis
@@
$\tau(w):=\begin{cases}a_{1}\dfrac{1}{\|\Xi(w)\|^{1-k_{1}}}&+a_{2}\dfrac{1}{\|\Xi(w)\|^{1-k_{2}}}+a_{3}\dfrac{1}{\|\Xi(w)\|}\\ &\mbox{ if }w\in\mathbb{R}^{d}\setminus{\rm Zer}(\Xi),\\ 0&\mbox{otherwise},\end{cases}$ (9)
with $a_{1},a_{2}>0,a_{3}\geq 0,k_{1}\in(0,1)$ and $k_{2}>1$.
@@
Let $h~:\mathbb{R}^{d}\longrightarrow\mathbb{R}^{d}$ be a Lipschitz continuous single-valued mapping with Lipschitz constant $\alpha$. Let $F:\mathbb{R}^{d}\longrightarrow\mathbb{R}^{d}$ be a $\beta$- Lipschitz continuous single-valued mapping. Let $g~:\Omega\longrightarrow\mathbb{R}\cup\{+\infty\}$ be a proper, convex, and lower semicontinuous function. Assume that Conditions in Assumption (A) hold.
Let $\overline{w}\in\mathbb{R}^{d}$ be the unique equilibrium point of the dynamical system (8). Then the solution $\overline{w}\in\mathbb{R}^{d}$ of the GIMVIP (1) is a fixed-time stable equilibrium point of (8) for any $k_{1}\in(0,1)$ and $k_{2}>1$ and the following time estimate holds:
@@
Let $V:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be the Lyapunov function defined by:
@@
$=$ $-\Big(a_{1}\|\Xi(w)\|^{k_{1}}+a_{2}\|\Xi(w)\|^{k_{2}}+a_{3}\Big)\langle w-\overline{w},\Xi(w)\rangle$
$\leq$ $-a_{1}\left(\rho-\Lambda\right)\dfrac{\|w-\overline{w}\|^{2}}{\|\Xi(w)\|^{1-k_{1}}}-a_{2}\big(\rho-\Lambda\big)\dfrac{\|w-\overline{w}\|^{2}}{\|\Xi(w)\|^{1-k_{2}}}-(\rho-\Lambda)a_{3}\dfrac{\|w-\overline{w}\|^{2}}{\|\Xi(w)\|}$
for all $w\in\mathbb{R}^{d}\setminus\{\overline{w}\}$.
Using Part (iii) of Lemma 7 one obtains that
$\dot{V}\leq$ $-\dfrac{a_{1}(\rho-\Lambda)}{\Gamma^{1-k_{1}}}\|w-\overline{w}\|^{1+k_{1}}-(\rho-\Lambda)^{k_{2}}\Big(a_{2}+a_{3}\|\Xi(w)\|^{k_{2}}\Big)\|w-\overline{w}\|^{1+k_{2}}$
$\leq$ $C_{1}(k_{1})\|w-\overline{w}\|^{1+k_{1}}-C_{2}(k_{2},u)\|w-\overline{w}\|^{1+k_{2}},$
where $C_{1}(k_{1})=\dfrac{a_{1}(\rho-\Lambda)}{\Gamma^{1-k_{1}}}$ and $C_{2}(k_{2},u)=(\rho-\Lambda)^{k_{2}}\Big(a_{2}+a_{3}\|\Xi(w)\|^{k_{2}}\Big)$. Since $\rho>\Lambda$ and $a_{i}>0,i=1,2,3$ it holds that $q_{1}>0,q_{2}>0$ for all $0<r_{1}<1,r_{2}>1$.
$\dot{V}\leq$ $C_{1}\left(k_{1}\right)\left(\|w-\overline{w}\|^{2}\right)^{\frac{1+k_{1}}{2}}-C_{2}(k_{2},u)\left(\|w-\overline{w}\|^{2}\right)^{\frac{1+k_{2}}{2}}$
$\leq$ $-\Big(A_{1}V(w)^{s_{1}}+A_{2}V(w)^{s_{2}}\Big),$ (10)
here $A_{i}=2^{s_{i}}C_{i}$ and $s_{i}=\frac{1+k_{i}}{2}$, for $i=1,2$. Note that $A_{1}>0,0.5<s_{1}<1$ for any $k_{1}\in(0,1)$ and $A_{2}>0,s_{2}>1$ for any $k_{2}>1$. Using Theorem 3, we conclude the conclusion of the theorem holds. ∎
@@
In line with Remark 3, the above result ensures fixed-time stability for IMVIPs when $h$ is the identity mapping, and for IVIPs when $g\equiv 0$. Moreover, if $F$ is the identity mapping, fixed-time stability is obtained for MVIPs, which reduce to classical VIPs when $g\equiv 0$.
When $F$ is the identity mapping, the resulting dynamical system (8)–(9) coincides with the one studied in [16]. Hence, as shown in [16], if $h\equiv\nabla f$ for some differentiable function $f$, the system (8)–(9) yields fixed-time stability for the constrained optimization problem

### v2

# Novel Dynamical Systems with Finite-Time and Predefined-Time Stability for Generalized Inverse Mixed Variational Inequality Problems

@@ Abstract
Two novel continuous-time dynamical systems are proposed to study the finite-time and predefined-time stability of solutions to GIMVIPs in finite-dimensional Hilbert spaces. Under suitable assumptions on the involved operators and model parameters, Lyapunov-based techniques are employed to establish finite-time and predefined-time convergence of the generated trajectories.
Although both dynamical systems exhibit accelerated convergence, the settling time of the finite-time stable system depends on the initial condition, whereas the predefined-time stable system admits a uniform upper bound on the convergence time that is independent of the initial state and can be explicitly prescribed through user-selected parameters. Moreover, by applying a forward Euler discretization to the continuous-time dynamics, a proximal point-type iterative algorithm is derived, and its fixed-time convergence property is rigorously analyzed. Numerical experiments are provided to illustrate the effectiveness and advantages of the proposed methods.
@@
Finite-time stability, predefined-time Stability, proximal point Method, Dynamical System, generalized Inverse mixed Variational Inequality Problem.
@@ 1 Introduction
Let $\langle\cdot,\cdot\rangle$ and $\|\cdot\|$ denote the inner product and the associated norm in $\mathbb{R}^{d}$, respectively. This paper is concerned with a class of inverse equilibrium problems known as the generalized inverse mixed variational inequality problem (GIMVIP). Given a nonempty closed and convex set $\Omega\subset\mathbb{R}^{d}$, the GIMVIP consists in finding a vector $\bar{w}\in\mathbb{R}^{d}$ such that
$F(\bar{w})\in\Omega\quad\text{and}\quad\langle h(\bar{w}),v-F(\bar{w})\rangle+g(v)-g(F(\bar{w}))\geq 0,\quad\forall v\in\Omega.$ (1)
@@
Problem (1) provides a unifying framework that encompasses several important models in variational analysis. If $h$ is the identity mapping, then (1) reduces to the inverse mixed variational inequality problem introduced in [18]. When $g\equiv 0$, it collapses to the inverse variational inequality problem (IVIP), originally proposed by He and Liu [13]. On the other hand, if $F$ is the identity operator, (1) becomes a mixed variational inequality problem (MVIP) and further reduces to the classical variational inequality problem (VIP) when $g\equiv 0$.
@@
Variational inequality problems constitute a fundamental modeling framework for equilibrium phenomena arising in optimization, saddle-point problems, Nash equilibria in noncooperative games, and fixed-point theory; see, for instance, [4, 5]. Beyond direct formulations, inverse variational inequality problems have received increasing attention due to their ability to infer unknown system parameters from observed equilibrium states. Such inverse models naturally arise in applications where equilibrium behavior is observable, whereas the underlying cost structures, control parameters, or governing operators are not directly accessible.
@@
In recent years, inverse variational inequality models have been successfully applied in traffic network analysis, economic equilibrium modeling, and telecommunication systems. To further enhance modeling flexibility, inverse mixed variational inequality problems were proposed by incorporating nonsmooth and composite terms, allowing richer representations of practical constraints. Subsequent developments led to generalized inverse mixed variational inequality formulations, which subsume many existing inverse models and provide a comprehensive framework for inverse equilibrium analysis. Motivated by these advances, the present work focuses on GIMVIPs as a versatile and unified inverse equilibrium model.
@@
Alongside model generalization, the design of efficient solution methods remains a central challenge. Continuous-time dynamical systems have emerged as a powerful tool for solving variational inequalities and related equilibrium problems. By reformulating equilibrium conditions as dynamical processes, one can exploit Lyapunov stability theory to analyze convergence properties. Compared with purely discrete-time schemes, continuous-time approaches offer conceptual simplicity, potential for real-time implementation, and a natural interpretation from a feedback control perspective.
@@
Most existing dynamical system approaches for variational and inverse variational inequality problems guarantee asymptotic or exponential stability. Specifically, dynamical system techniques have been applied to optimization problems [2, 6, 11, 12, 19, 23], sparse optimization [10, 12, 28], control and multi-agent systems [29], mixed variational inequality problems [9, 14], as well as variational inequality and equilibrium problems [15, 16]. More recently, Nam et al. [26] investigated finite-time stability for generalized monotone inclusion problems, while fixed-time stable dynamical systems for such inclusions were developed in [27]. In addition, Ju et al. [17] employed continuous-time dynamical systems to establish fixed-time stability for pseudo-monotone mixed variational inequality problems.
@@
For inverse and inverse quasi-variational inequality problems, dynamical system methods have also been explored; see, for example, [30]. In particular, Dey and Reich [8] proposed a first-order continuous-time dynamical system for inverse quasi-variational inequality problems and established global exponential stability. Subsequently, Vuong and Thanh [25] analyzed a time-discretized version of a related system and obtained linear convergence results under a more general moving set condition. More recently, Tangkhawiwetkul [24] studied generalized inverse mixed variational inequality problems using dynamical system techniques and derived asymptotic as well as exponential stability results.
@@
However, in many applications requiring rapid, predictable, and time-critical convergence, asymptotic or exponential stability may be insufficient. Finite-time stability, introduced in [3], guarantees convergence within a finite settling time, although this time typically depends on the initial condition. To address this limitation, Polyakov [22] introduced the concept of fixed-time stability, for which the convergence time admits a uniform upper bound independent of the initial state. Despite their practical relevance, finite-time and fixed-time convergence properties have not yet been systematically investigated for generalized inverse mixed variational inequality problems.
Beyond finite-time and fixed-time stability, an even stronger notion, known as predefined-time stability, was introduced by Sánchez-Torres et al. [Sanchez2014]. In contrast to fixed-time stability, where convergence is guaranteed within an unknown but bounded time, predefined-time stability allows the convergence time to be explicitly prescribed in advance through user-selected parameters and remains independent of the initial condition. This feature provides enhanced flexibility and autonomy in system design, enabling practitioners to enforce convergence within a desired time horizon that meets practical control or performance requirements.
Predefined-time stability has been actively studied in control and synchronization contexts. For instance, Anguiano-Gijón et al. [1] developed Lyapunov-function-based control schemes for predefined-time synchronization of chaotic systems, while Liu et al. [20] proposed novel predefined-time stability lemmas and applied them to synchronization problems. These studies demonstrate the effectiveness of predefined-time stability in facilitating coordination and information exchange among interconnected systems. Nevertheless, the incorporation of predefined-time stability into optimization and inverse equilibrium problem frameworks remains largely unexplored.
@@
Motivated by the above observations, this paper develops novel first-order continuous-time dynamical systems for solving generalized inverse mixed variational inequality problems and establishes their finite-time, fixed-time, and predefined-time stability properties. By applying a forward Euler discretization, we further derive a proximal point-type iterative algorithm and show that the generated sequence converges to the unique solution of the GIMVIP within a prescribed time horizon. The proposed framework not only generalizes existing dynamical system approaches for inverse variational inequality problems but also provides efficient and user-controllable solution methods for generalized inverse equilibrium models.
The remainder of the paper is organized as follows. Section 2 introduces preliminary material, including notions of finite-time, fixed-time, and predefined-time stability, and establishes the correspondence between GIMVIPs and the equilibrium points of an associated proximal dynamical system. Section 3 investigates the finite-time stability of a newly proposed dynamical system. Section 4 is devoted to the fixed-time and predefined-time stability analysis of an alternative system. Section 5 studies the consistency of its time discretization, ensuring convergence within a prescribed number of iterations. Numerical experiments illustrating the effectiveness of the proposed methods are reported in Section 6. Finally, Section 7 concludes the paper with a summary of the main results.
@@ 2.2 Equilibrium points and stability
If we can adjust the system parameters such that the settling time in a fixed-time stable dynamical system can be predicted, then the equilibrium is called a predefined-time stable equilibrium.
@@ 2.3 A nominal dynamical system
    $\langle w-\overline{w},\Xi(w)\rangle\geq(\rho-\Lambda)\|w-\overline{w}\|^{2}>0\ \forall w\in\mathbb{R}^{d},\ w\neq\overline{w}$.
@@ 3 Finite-time stability analysis
4 Fixed-time and predefined-time stability analysis
@@
$\tau(w):=\begin{cases}a_{1}\dfrac{1}{\|\Xi(w)\|^{1-k_{1}}}&+a_{2}\dfrac{1}{\|\Xi(w)\|^{1-k_{2}}}+a_{3}\dfrac{1}{\|\Xi(w)\|^{k_{3}}}\\ &\mbox{ if }w\in\mathbb{R}^{d}\setminus{\rm Zer}(\Xi),\\ 0&\mbox{otherwise},\end{cases}$ (9)
with $a_{1},a_{2}>0,a_{3}>0,k_{1}\in(0,1)$, $k_{3}\geq 0$ and $k_{2}>1$.
Remark 4.
When $F$ is the identity mapping and $k_{3}=0$, the dynamical system (8) reduces to the system proposed in [31], which was used to establish predefined-time stability for MVIPs. Likewise, when $F$ is the identity mapping and $k_{3}=1$, the dynamical system (8) coincides with the one studied in [17], where fixed-time stability for MVIPs was established.
@@
The following proposition provides a sufficient condition for the existence of a solution to the system (8).
Proposition 13.
Consider the dynamical system (8). If $\Xi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ be a locally Lipschitz continuous vector field such that
[…]
then the right-hand side of (8) is continuous for all $w\in\mathbb{R}^{n}$, and for any given initial condition, the solution of (8) exists and is uniquely determined for all $t\geq 0$.
Proof.
Using the same arguments as in the proof of Lemma 3 in [31], we obtain the desired conclusion.. ∎
The following lemma from [31] provides sufficient conditions for predefined-time stability.
Lemma 14.
[31] Consider the system (2). If there exists a positive definite unbounded radial function $V:\mathbb{R}^{d}\rightarrow\mathbb{R}$, and a user-defined parameter $T_{d}$ such that the following conditions hold:
(i) $V(w(t))=0\Leftrightarrow\dot{w}(t)=0$.
(ii) For any $V(w(t))>0$, there exist $c_{1},c_{2},c_{3},r_{1},r_{2},T_{d},G_{d}>0$, where $r_{2}>1$, and $0<r_{1}<1$, such that the following inequality holds:
$\dot{V}(w(t))\leq-\frac{G_{d}}{T_{d}}\left(c_{1}V^{r_{1}}(w(t))+c_{2}V^{r_{2}}(w(t))+c_{3}V(w(t))\right),$ (11)
then the system (2) reaches to fixed-time stability within the predefined time $T_{d}$, where
$G_{d}=\frac{1}{c_{3}(1-r_{1})}\ln\left(1+\frac{c_{3}}{c_{1}}\right)+\frac{1}{c_{3}(r_{2}-1)}\ln\left(1+\frac{c_{3}}{c_{2}}\right)$ (12)
@@
Let $h~:\mathbb{R}^{d}\longrightarrow\mathbb{R}^{d}$ be a Lipschitz continuous single-valued mapping with Lipschitz constant $\alpha$. Let $F:\mathbb{R}^{d}\longrightarrow\mathbb{R}^{d}$ be a single-valued mapping. Let $g~:\Omega\longrightarrow\mathbb{R}\cup\{+\infty\}$ be a proper, convex, and lower semicontinuous function. Assume that Conditions in Assumption (A) hold.
Let $\overline{w}\in\mathbb{R}^{d}$ be the unique equilibrium point of the dynamical system (8). Then
For $l=1$, the solution $\overline{w}\in\mathbb{R}^{d}$ of the GIMVIP (1) is a fixed-time stable equilibrium point of (8) for any $k_{1}\in(0,1)$ and $k_{2}>1$ and the following time estimate holds:
@@
In particular, if $k_{3}=0$, the (8) can reaches the equilibrium point within a predefined time $T_{d}$, where $T_{d}$ is a user-defined parameter, and $G_{d}=\frac{1}{a_{3}(\rho-\Lambda)\left(1-k_{1}\right)}\ln\left(1+\frac{a_{3}\sqrt{2}^{1-k_{1}}\Gamma^{1-k_{1}}}{a_{1}}\right)+\frac{1}{a_{3}(\rho-\Lambda)\left(k_{2}-1\right)}\ln\left(1+\frac{a_{3}\sqrt{2}^{1-k_{2}}(\rho-\Lambda)^{1-k_{2}}}{a_{2}}\right)$.
@@
By Theorem 6 the generalized inverse variational inequality problem has a unique solution. By Part (iv) of Lemma (3), all assumptions in Proposition 13 are satisfied. Hence, the system (8) has a unique solution. Consider the following Lyapunov function
$V(w)=\frac{1}{2}\left\|w-\overline{w}\right\|^{2}$. Let $V:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be the Lyapunov function defined by:
@@
$=$ $-\frac{G_{d}}{T_{d}}\Big(a_{1}\|\Xi(w)\|^{k_{1}}+a_{2}\|\Xi(w)\|^{k_{2}}+a_{3}\Big)\langle w-\overline{w},\Xi(w)\rangle$
$\leq$ $-\frac{G_{d}}{T_{d}}\left(a_{1}\left(\rho-\Lambda\right)\dfrac{\|w-\overline{w}\|^{2}}{\|\Xi(w)\|^{1-k_{1}}}-a_{2}\big(\rho-\Lambda\big)\dfrac{\|w-\overline{w}\|^{2}}{\|\Xi(w)\|^{1-k_{2}}}-(\rho-\Lambda)a_{3}\dfrac{\|w-\overline{w}\|^{2}}{\|\Xi(w)\|^{k_{3}}}\right)$ (13)
for all $w\in\mathbb{R}^{d}\setminus\{\overline{w}\}$. Using Part (iii) of Lemma 7 one obtains that
$\dot{V}\leq$ $\frac{G_{d}}{T_{d}}\left(-\dfrac{a_{1}(\rho-\Lambda)}{\Gamma^{1-k_{1}}}\|w-\overline{w}\|^{1+k_{1}}-(\rho-\Lambda)^{k_{2}}\Big(a_{2}+a_{3}\|\Xi(w)\|^{1-k_{2}-k_{3}}\Big)\|w-\overline{w}\|^{1+k_{2}}\right)$
$\leq$ $-C_{1}(k_{1})\|w-\overline{w}\|^{1+k_{1}}-C_{2}(k_{2},u)\|w-\overline{w}\|^{1+k_{2}},$ (14)
where $C_{1}(k_{1})=\frac{G_{d}}{T_{d}}\cdot\dfrac{a_{1}(\rho-\Lambda)}{\Gamma^{1-k_{1}}}$ and $C_{2}(k_{2},u)=\frac{G_{d}}{T_{d}}(\rho-\Lambda)^{k_{2}}\Big(a_{2}+a_{3}\|\Xi(w)\|^{1-k_{2}-k_{3}}\Big)$. Since $\rho>\Lambda$ and $a_{i}>0,i=1,2,3$ it holds that $C_{1}(k_{1})>0,C_{2}(k_{2},u)>0$ for all $0<r_{1}<1,r_{2}>1$. By setting
$A_{i}=2^{\frac{1+k_{1}}{2}}\frac{G_{d}}{T_{d}}\cdot\dfrac{a_{1}(\rho-\Lambda)}{\Gamma^{1-k_{1}}}$
and choosing
$A_{2}\in\left[2^{\frac{1+k_{2}}{2}}\frac{G_{d}}{T_{d}}(\rho-\Lambda)^{k_{2}}a_{2},2^{\frac{1+k_{2}}{2}}\frac{G_{d}}{T_{d}}(\rho-\Lambda)^{k_{2}}\Big(a_{2}+a_{3}\|\Xi(w(0))\|^{1-k_{2}-k_{3}}\Big)\right]$
we derive from (4) that
$\dot{V}\leq$ $-\Big(A_{1}V(w)^{s_{1}}+A_{2}V(w)^{s_{2}}\Big),$ (15)
here $s_{i}=\frac{1+k_{i}}{2}$, for $i=1,2$. Note that $A_{1}>0,0.5<s_{1}<1$ for any $k_{1}\in(0,1)$ and $A_{2}>0,s_{2}>1$ for any $k_{2}>1$. Using Theorem 3, we conclude the conclusion of the theorem holds.
When $k_{3}=0$, from (4) we have
$\dot{V}$ $=\left\langle w-\overline{w},\dot{w}\right\rangle$
$=-\frac{G_{d}}{T_{d}}\left\langle w-\overline{w},a_{1}\frac{\Xi(w)}{\|\Xi(w)\|^{1-k_{1}}}+a_{2}\frac{\Xi(w)}{\|\Xi(w)\|^{1-k_{2}}}+a_{3}\Xi(w)\right\rangle$
$\leq\frac{G_{d}}{T_{d}}\left(-a_{1}(\rho-\Lambda)\frac{\left\|w-\overline{w}\right\|^{2}}{\|\Xi(w)\|^{1-k_{1}}}-a_{2}(\rho-\Lambda)\frac{\left\|w-\overline{w}\right\|^{2}}{\|\Xi(w)\|^{1-k_{2}}}-a_{3}(\rho-\Lambda)\left\|w-\overline{w}\right\|^{2}\right)$
$\leq\frac{G_{d}}{T_{d}}\left(-\frac{a_{1}(\rho-\Lambda)}{\Gamma^{1-k_{1}}}\left\|w-\overline{w}\right\|^{1+k_{1}}-a_{2}(\rho-\Lambda)\frac{\left\|w-\overline{w}\right\|^{2}}{\|\Xi(w)\|^{1-k_{2}}}-a_{3}(\rho-\Lambda)\left\|w-\overline{w}\right\|^{2}\right)$
$\leq\frac{G_{d}}{T_{d}}\left(-\frac{a_{1}(\rho-\Lambda)}{\Gamma^{1-k_{1}}}\left\|w-\overline{w}\right\|^{1+k_{1}}-a_{2}(\rho-\Lambda)^{k_{2}}\left\|w-\overline{w}\right\|^{1+k_{2}}-a_{3}(\rho-\Lambda)\left\|w-\overline{w}\right\|^{2}\right)$
$=\frac{G_{d}}{T_{d}}\left(-\frac{a_{1}(\rho-\Lambda)}{\Gamma^{1-k_{1}}}(2V)^{\frac{1+k_{1}}{2}}-a_{2}(\rho-\Lambda)^{k_{2}}(2V)^{\frac{1+k_{2}}{2}}-2a_{3}(\rho-\Lambda)V\right)$
$=\frac{G_{d}}{T_{d}}\left(-\left(\sqrt{2}^{1+k_{1}}\frac{a_{1}(\rho-\Lambda)}{\Gamma^{1-k_{1}}}V^{\frac{1+k_{1}}{2}}+\sqrt{2}^{1+k_{2}}a_{2}(\rho-\Lambda)^{k_{2}}V^{\frac{1+k_{2}}{2}}+2a_{3}(\rho-\Lambda)V\right)\right)$
$=\frac{G_{d}}{T_{d}}\left(-\tau_{1}V^{\frac{1+k_{1}}{2}}-\tau_{2}V^{\frac{1+k_{2}}{2}}-\tau_{3}V\right)$ (16)
where $\tau_{1}=\sqrt{2}^{1+k_{1}}\frac{a_{1}(\rho-\Lambda)}{\Gamma^{1-k_{1}}}>0,\tau_{2}=\sqrt{2}^{1+k_{2}}a_{2}(\rho-\Lambda)^{k_{2}}>0,\tau_{3}=2a_{3}(\rho-\Lambda)$. Take
$G_{d}=$ $\frac{2}{\tau_{3}\left(1-k_{1}\right)}\ln\left(1+\frac{\tau_{3}}{\tau_{1}}\right)+\frac{2}{\tau_{3}\left(k_{2}-1\right)}\ln\left(1+\frac{\tau_{3}}{\tau_{2}}\right)$
$=$ $\frac{1}{a_{3}(\rho-\Lambda)\left(1-k_{1}\right)}\ln\left(1+\frac{a_{3}\sqrt{2}^{1-k_{1}}\Gamma^{1-k_{1}}}{a_{1}}\right)$
$+\frac{1}{a_{3}(\rho-\Lambda)\left(k_{2}-1\right)}\ln\left(1+\frac{a_{3}\sqrt{2}^{1-k_{2}}(\rho-\Lambda)^{1-k_{2}}}{a_{2}}\right).$
Then, by (4) and Lemma 14, we conclude that the equilibrium point of the dynamical system (8) is reached within the predefined time $T_{d}$. Moreover, for any initial condition $w(0)\in\mathbb{R}^{n}$, the settling time satisfies
$T(w(0))\leq T_{\max}=T_{d}.$
The proof is completed. ∎
@@
In line with Remark 3, when $l=1$ the above result ensures fixed-time stability for IMVIPs when $h$ is the identity mapping, and for IVIPs when $g\equiv 0$. Moreover, if $F$ is the identity mapping, fixed-time stability is obtained for MVIPs, which reduce to classical VIPs when $g\equiv 0$. When $l=0$, the above result implies the predefined stability for an equilibrium point of corresponding problems with a particular choice of mapping $F,h$ and the function $g$.
When $k_{3}=1$ and $F$ is the identity mapping, the resulting dynamical system (8) coincides with the one studied in [17]. Hence, as shown in [17], if $h\equiv\nabla f$ for some differentiable function $f$, the system (8) yields fixed-time stability for the constrained optimization problem